How do I Graph X + Y > 1?
I have been trying for two days! Please show me.

Answer by richard1234(7193)

Two days?

All you need to do is graph x+y = 1, then shade the area above it, since you want x+y to be greater than 1. To graph it more easily, write it as y > 1 - x, graph the equation y = 1 - x, then shade everything above it. It is sometimes conventional to draw a dashed line through y = 1 - x, indicating that it is not part of the function.